# A Wiener process perspective on local intrinsic dimension estimation methods

### v1

@@ Abstract
Local intrinsic dimension (LID) estimation methods have received a lot of attention in recent years thanks to the progress in deep neural networks and generative modeling. In opposition to old non-parametric methods, new methods use generative models to approximate diffused dataset density and scale the methods to high-dimensional datasets like images. In this paper, we investigate the recent state-of-the-art parametric LID estimation methods from the perspective of the Wiener process. We explore how these methods behave when their assumptions are not met. We give an extended mathematical description of those methods and their error as a function of the probability density of the data.
@@ 1 Introduction
LID estimation has gained increasing attention in recent years as part of the fast-growing field of topological data analysis. It was able to progress from non-parametric methods to parametric ones thanks to the latest progress in the field of generative modeling. LID estimation methods are algorithms for estimating the manifold’s dimension from which the data point $x$ was sampled. In those methods, we assume that the data lie on the union of one or more manifolds that may be of different dimensions.
@@
The estimation of intrinsic dimension is a critical issue in data analysis and machine learning [Ansuini et al., 2019, Li et al., 2018, Rubenstein et al., 2018] and was investigated in relation to dimensionality reduction and clustering [Vapnik, 2013, Kleindessner and Luxburg, 2015, Camastra and Staiano, 2016], analyzing the training and representation learning processes within deep neural networks [Li et al., 2018, Ansuini et al., 2019, Pope et al., 2020, Loaiza-Ganem et al., 2024], verifying the union of manifolds hypothesis for images Brown et al. [2022], and used to improve out-of-distribution detection algorithm Kamkari et al. [2024a].
@@
Before Tempczyk et al. [2022], two approaches were presented to solve the intrinsic dimension estimation problem, both based on non-parametric methods. The first approach employs global methods such as PCA and its variants; see, e.g., Fukunaga and Olsen [1971], Minka [2000], Fan et al. [2010]. These methods are known to suffer from issues related to a manifold curvature and non-uniformity of the data distribution. They also assumed that data lies on a single manifold of constant dimension, so dimensionality is the same for all $x$. The second approach is based on local methods that explore the geometric properties of neighborhoods Johnsson et al. [2014], Levina and Bickel [2004] calculating some statistics using points from the neighborhood of $x$. Although all aforementioned methods perform reasonably well for a small number of dimensions, the higher dimensionality affects their performance Tempczyk et al. [2022], Campadelli et al. [2015], Camastra and Staiano [2016].
@@
Very recently, a new approach emerged for the local methods, which is based on a two-step procedure. In the first step, the dataset is perturbed with some noise, and in the second step, its dimensionality is estimated using various techniques. These include generative models to analyze changes in density Tempczyk et al. [2022], Kamkari et al. [2024a] or in singular values of the Jacobian  Horvat and Pfister [2022, 2024] for different noise magnitudes, and analyzing rank of scores from diffusion model as mentioned in Stanczuk et al. [2024].
@@
Although the new algorithms provide state-of-the-art performance for large and real-life data sets, it is crucial for their performance that adding noise to the data set is an injective process, which means that the resulting densities after adding noise uniquely identify the original distribution of the data set. Otherwise, it is impossible to neither uniquely reverse the process nor analyze the original structure of the data set. One example in the theoretical model that does not admit such a problem is a flat manifold with a uniform distribution of data points. In such a case those algorithms perform best, as showed in experimental results. In other cases, a hypothetical line of attack would be to decrease the magnitude of noise added to the data set so that the manifold is approximately flat and uniform in the scale of the noise magnitude. This, however, is not possible in practice for several reasons. One is that data is often a set of discrete points (especially in the case of audio and visual data), and considering the noise of much smaller magnitude than the minimum distance between points does not lead to any meaningful process. Moreover, neural net training is done with finite precision and the stochastic gradient descent procedure introduces noise to the density estimates, so very small changes in density cannot be observed in practice, which may lead to poor quality estimates of LID.
@@
In this paper, we point out and exploit the fact that adding Gaussian noise of varying magnitudes can be seen as studying the evolution of the Wiener process describing the diffusion of particles (points of the dataset) in space. This point of view enables us to employ Fick’s Second Law of Diffusion to eliminate time derivatives from mathematical descriptions of state-of-the-art LID algorithms Tempczyk et al. [2022], Kamkari et al. [2024a], and replace them with spatial derivatives. Such considerations can be taken into account in the second step of the considered algorithms, leading to more accurate results.
Contribution
@@
    We explore the first step of existing algorithms in the language of Wiener process and calculate important cases of diffusion from lower-dimensional manifolds with non-uniform probability density into higher dimensional ambient space.
@@
    We recognize and define new categories (isolated and holistic algorithms) for the Wiener process based parametric LID estimation algorithm family and categorize existing algorithms accordingly.
@@
    We derive closed-form expressions for crucial parameters used in two state-of-the-art isolated LID estimation algorithms as a function of on-manifold density and manifold dimensionality, which can be viewed as closed-form expression for deviation from flat manifold with uniform distribution case for them.
@@ 2 Related work
The review of non-parametric methods for local and global intrinsic dimension estimation can be found in Campadelli et al. [2015], Camastra and Staiano [2016], and a broad comparison of those methods on bigger datasets (in terms of dimensionality) can be found in Tempczyk et al. [2022].
Although we do not analyze non-parametric methods in this paper, the interesting and worth mentioning recent work on non-parametric LID estimation is Bailey et al. [2022]. Authors explore the connection of LID to other well-known measures for complexity: entropy and statistical divergences and develop new analytical expressions for these quantities in terms of LID. The same authors in their follow-up paper Bailey et al. [2023] establish the relationships for cumulative Shannon entropy, entropy power, Bregman formulation of cumulative Kullback–Leibler divergence, and generalized Tsallis entropy variants, and propose four new estimators of LID, based on nearest neighbor distances.
2.1 Parametric methods for LID estimation
During the last few years many parametric methods for estimating LID emerged. In Zheng et al. [2022] the authors prove that VAE are capable of recovering the correct manifold dimension and demonstrate methods to learn manifolds of varying dimension across the data sample. Authors of Yeats et al. [2023] connects the adversarial vulnerability of score models with the geometry of the underlying manifold they capture. They show that minimizing the Dirichlet energy of learned score maps boosts their robustness while revealing LID of the underlying data manifold.
Wiener process based algorithms
Among parametric methods, there is a group of algorithms, that have one thing in common: they simulate a Wiener process on a dataset and directly use some properties of time evolving density to estimate LID. We can divide those algorithms into three groups:
@@ 2.1 Parametric methods for LID estimation
    LIDL Tempczyk et al. [2022] and its efficient and most accurate implementation using diffusion models called FLIPD Kamkari et al. [2024b]. These algorithms use the rate of change of the probability density at point $x$ during the Wiener process to estimate LID at $x$. For small diffusion times $t$ the logarithm of a density is a linear function of a logarithm of $t$, and the proportionality coefficient is equal $d-D$ for small $t$, where $d$ is manifold density and $D$ is ambient space dimensionality. Authors of FLIPD show, that their algorithm is efficient and our experiments show, that is more scalable than algorithms from the second group (and with high probability more accurate, but there is a lack of comparison between FLIPD and methods from second group) and more accurate than algorithms from the third group, which led us to the conclusion that at this moment this family of algorithms is state-of-the-art in LID estimation.
@@
    ID-NF Horvat and Pfister [2022] (using normalizing flows), and the follow-up paper Horvat and Pfister [2024] ID-DM (using diffusion models) algorithms analyze how a singular values of a Jacobian of a function transforming a standard normal distribution into a diffused dataset at $x$ evolves during the Wiener process. Authors use an observation, that when transforming a $d$-dimensional manifold into a $D$-dimensional Gaussian we have to expand space more in the off-manifold directions, especially for small diffusion times $t$.
@@
    The third group, for now, is composed of one algorithm: Stanczuk et al. [2024]. We will call this algorithm NB – Normal Bundle – following the convention from Kamkari et al. [2024b]. Authors of NB make an observation, that for small diffusion times $t$ the score function from the Score-based diffusion model Song et al. [2020] lies in the normal space to the manifold at $x$. The procedure is the following: calculate the score function for a bunch of diffused versions of $x$ for a fixed diffusion time $t$, stack them as column vectors into a matrix $S$, and estimate the rank of $S$, which is then used as an estimate of $D-d$.
@@ 3 Isolated and holistic algorithms
In this work are going to take a closer look at the Wiener process based algorithms described in the last paragraph of Section 2. Although all of those algorithms apply a Wiener process to the dataset during their first phase, when looking at their second phase, we can divide them into two groups: isolated (LIDL, FLIPD, NB) and holistic (ID-NF, ID-DM). The intermediate results of the first group used to calculate LID use only (assuming that a generative model approximates diffused data distribution $\rho_{t}$ perfectly) information about the shape of the data probability density function (without normalization constant). Their estimate depends on the gradient of the logarithm of $\rho_{t}$ around $x$ (NB) or $d\log\rho_{t}/d\log t$ at $x$ (LIDL, FLIPD). Because in practice $\rho_{t}$ is a function of the values of original data distribution $p_{S}$ in a ball of radius $r\approx 4\sqrt{t}$ around $x$ in the data space. What the values of $p_{S}$ outside this ball are do not matter in practice for isolated algorithms because the diffused particles in the Wiener process can travel longer distances than $r$ with very low probability (less than $3.7\cdot 10^{-5}$). When we add some new data points to the dataset far away from $x$, it doesn’t change the shape of $\psi$ close to $x$ (this operation only changes a normalization constant, which becomes 0 after taking a logarithm of $\psi$ and taking a derivative).
The holistic algorithms work quite differently. In the case of ID-NF and ID-DM, they calculate singular values of the Jacobian of the function $\zeta:\mathrm{R}^{N}\rightarrow\mathrm{R}^{N}$ transforming $\rho_{t}$ into a standard normal distribution. This $\zeta$ function strongly depends on the entire shape of $\rho_{t}$. When we change $\rho_{t}$ far away from $x$ we don’t change the shape of $\rho_{t}$ close to $x$. The same is not true for $\zeta(x)$. When we add many data points to the dataset far away from $x$ and we keep the latent distribution fixed we have to change the way we compress and stretch the space to match the new distribution. This property makes the analysis of the behavior of ID-NF and ID-DM much harder (maybe even impossible) as a function of density evolution around $x$.
To give an illustrative example of this behavior let’s imagine that our dataset is one-dimensional and consists of 10K points sampled from $\mathcal{N}\left(0,1\right)$. Because we are transforming it into $\mathcal{N}\left(0,1\right)$, $\zeta$ is an identity function. Now let’s add to the dataset another 10K points sampled from $\mathcal{N}\left(100,1\right)$. Now we have to stretch our space in some areas to transform one density into another and our $\zeta$ function (and its Jacobian) is far from being an identity function.
This approach can be extended for new algorithms in the future, and similar analysis can be carried out for the new algorithms from isolated group if only some intermediate results used to calculate LID can be expressed in the analytical form.
@@ 4 Wiener process perspective on LID estimation
In this section we present a new perspective on perturbing datasets, unifying the approaches seen in the algorithms presented in Tempczyk et al. [2022], Stanczuk et al. [2024], Horvat and Pfister [2022, 2024], Kamkari et al. [2024b]. As already mentioned, all these algorithms consist of two stages, the first of which amounts to perturbing the dataset with normally distributed random noise of fixed variance $t$. In the second stage, each of the algorithms utilizes the behavior of the perturbed density in the neighborhood of a fixed point under changes of the noise variance.
The first phase of each algorithm can be interpreted as applying the Wiener process to the points in the dataset. Afterward, the resulting set of points is used to train a generative model (depending on the chosen algorithm) to estimate the distribution of the dataset undergoing the Wiener process at time $t$. From the point of view of differential equations, the distribution density function of the diffused dataset is described by Fick’s Second Law of Diffusion.
@@
Now, let us get to the specifics. Given a dataset embedded in $\mathbb{R}^{D}$, we assume that it has been drawn from some latent union of submanifolds $S$ endowed with a probability measure $p_{S}$ (which can be naturally treated as probability measure on $\mathbb{R}^{D}$). The goal of Local Intrinsic Dimension estimation is to find out the dimension of $S$ at any point of the dataset.
@@
This is the density of normal distribution on $\mathbb{R}^{D}$ with covariance matrix $tI$. It is the fundamental solution of the differential equation given by Fick’s second Law of Diffusion (1) with $C=1/2$. In our context, this means that the convolution
$\rho_{t}=p_{S}*\phi_{t}$ (3)
@@
To limit the complexity introduced by curvature, we will consider from now on only flat manifolds. This means that, without loss of generality, we may assume that $S$ is the first factor in product decomposition $\mathbb{R}^{D}=\mathbb{R}^{d}\times\mathbb{R}^{D-d}$. We will denote the coordinates of $\mathbb{R}^{d}$ and $\mathbb{R}^{D-d}$ by $x$ and $y$, respectively. If we moreover assume that $p_{S}$, now a probability distribution on $\mathbb{R}^{d}$, has a density $\psi\colon\mathbb{R}^{d}\to\mathbb{R}$, we can separate variables in (3), obtaining
$\rho_{t}(x,y)=(\psi*\phi_{t}^{d})(x)\phi_{t}^{D-d}(y)$ (4)
for $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{D-d}$.
The decomposition $\mathbb{R}^{D}=\mathbb{R}^{d}\times\mathbb{R}^{D-d}$ gives rise to the decomposition of the Laplacian on $\mathbb{R}^{D}$ into Laplacians on the factors, namely $\Delta=\Delta_{x}+\Delta_{y}$, by simply grouping the derivatives with respect to $x_{i}$ and $y_{i}$ coordinates.
Lemma 4.1.

### v2

@@ Abstract
Local intrinsic dimension (LID) estimation methods have received a lot of attention in recent years thanks to the progress in deep neural networks and generative modeling. In opposition to old non-parametric methods, new methods use generative models to approximate diffused dataset density to scale the methods to high-dimensional datasets (e.g. images). In this paper, we investigate the recent state-of-the-art parametric LID estimation methods from the perspective of the Wiener process. We explore how these methods behave when their assumptions are not met. We give an extended mathematical description of those methods and their error as a function of the probability density of the data.
@@ 1 Introduction
LID estimation has gained increasing attention in recent years as part of the fast-growing field of topological data analysis. It was able to progress from non-parametric to parametric methods thanks to the latest progress in the field of generative modeling. LID estimation methods are algorithms for estimating the manifold’s dimension from which the data point $x$ was sampled. In these methods, we assume that the data lie on the union of one or more manifolds that may be of different dimensions.
@@
The estimation of intrinsic dimension is substantial for data analysis and machine learning (Ansuini et al. 2019; Li et al. 2018; Rubenstein, Schoelkopf, and Tolstikhin 2018) and was investigated in relation to dimensionality reduction and clustering (Vapnik 2013; Kleindessner and Luxburg 2015; Camastra and Staiano 2016), analyzing the training and representation learning processes within deep neural networks (Li et al. 2018; Ansuini et al. 2019; Pope et al. 2020; Loaiza-Ganem et al. 2024), verifying the union of manifolds hypothesis for images (Brown et al. 2022), and used to improve out-of-distribution detection algorithm (Kamkari et al. 2024a).
@@
Prior to introducing LIDL (Tempczyk et al. 2022), two approaches to solve the intrinsic dimension estimation problem were presented. The first employs global methods; see, e.g., Fukunaga and Olsen (1971); Minka (2000); Fan et al. (2010). These methods are known to suffer from issues related to a manifold curvature and non-uniformity of the data distribution. They also assume that data lies on a single manifold of constant dimension, so dimensionality is the same for all $x$. The second approach is based on local non-parametric methods that explore the geometric properties of neighborhoods (Johnsson, Soneson, and Fontes 2014; Levina and Bickel 2004) calculating some statistics using points from the neighborhood of $x$. Although all aforementioned methods perform reasonably well for a small number of dimensions, the higher dimensionality negatively affects their performance (Tempczyk et al. 2022; Campadelli et al. 2015; Camastra and Staiano 2016).
@@
Very recently, a new approach emerged for the local methods, which is based on a two-step procedure. In the first step, the dataset is perturbed with some noise, and in the second step, its dimensionality is estimated using various techniques. These include generative models to analyze changes in density (Tempczyk et al. 2022; Kamkari et al. 2024b) or in singular values of the Jacobian  (Horvat and Pfister 2022, 2024) for different noise magnitudes, and analyzing rank of scores from diffusion model as recently presented by Stanczuk et al. (2024).
@@
While new algorithms provide state-of-the-art results for large, real-life datasets, it is vital for their performance that adding noise to the dataset should be as close as possible to an injective process. This means that the resulting densities after adding noise should uniquely identify the original distribution of the dataset. Otherwise, it is impossible to neither uniquely reverse the process nor analyze the original structure of the dataset. One example in the theoretical model that does not admit such a problem is a flat manifold with a uniform distribution of data points. In such cases these algorithms perform best, as shown in experimental results.
In other cases, a hypothetical line of research would be to decrease the magnitude of noise added to the data set so that the manifold is approximately flat and has locally uniform data density in the scale of the noise magnitude. This, however, is not possible in practice for several reasons. For instance, data is often a set of discrete points (especially in audio and visual modality), and considering the noise of magnitude much smaller than the minimum distance between points does not lead to any meaningful process. Moreover, neural net training is done with finite precision and the stochastic gradient descent procedure introduces noise to the density estimates, so very small changes in density cannot be observed in practice, which may lead to poor quality estimates of LID.
@@
In this paper, we point out and exploit the fact that adding Gaussian noise of varying magnitudes can be seen as studying the evolution of the Wiener process describing the diffusion of particles (points of the dataset) in the ambient space. This point of view enables us to employ Fick’s Second Law of Diffusion to eliminate time derivatives from mathematical descriptions of state-of-the-art LID algorithms (Tempczyk et al. 2022; Kamkari et al. 2024b), and replace them with spatial derivatives. Such considerations can be taken into account in the second step of the considered algorithms, leading to more accurate results.
Contribution.
@@
    We recognize and define new categories (isolated and holistic algorithms) for the Wiener process-based parametric LID estimation algorithm family and categorize the existing algorithms accordingly.
@@
    We explore the first step of existing algorithms in the language of Wiener processes and calculate important cases of diffusion from lower-dimensional manifolds with non-uniform probability density into ambient space.
@@
    We derive closed-form expressions for important parameters used in two state-of-the-art isolated LID estimation algorithms as a function of on-manifold density and manifold dimensionality, which can be viewed as closed-form expressions for deviation from a flat manifold with uniform distribution case.
@@ 2 Related work
The review of the non-parametric methods for local and global intrinsic dimension estimation can be found in the work of Campadelli et al. (2015), or Camastra and Staiano (2016). Tempczyk et al. (2022) compared these methods on bigger datasets in terms of their dimensionality.
Although we do not analyze non-parametric methods in this paper, it is worth mentioning a recent work on non-parametric LID estimation, in which Bailey, Houle, and Ma (2022) explore the connection of LID to other well-known measures for complexity: entropy and statistical divergences, and develop new analytical expressions for these quantities in terms of LID. Consequently, Bailey, Houle, and Ma (2023) establish the relationships for cumulative Shannon entropy, entropy power, Bregman formulation of cumulative Kullback-Leibler divergence, and generalized Tsallis entropy variants, and propose four new estimators of LID, based on nearest neighbor distances.
During the last few years many parametric methods for estimating LID emerged. Zheng et al. (2022) prove that VAE are capable of recovering the correct manifold dimension and demonstrate methods to learn manifolds of varying dimensions across the data sample. Yeats et al. (2023) connect the adversarial vulnerability of score models with the geometry of the underlying manifold they capture. They show that minimizing the Dirichlet energy of learned score maps boosts their robustness while revealing the LID of the underlying data manifold.
Wiener process-based algorithms.
Regarding parametric methods, there is a group of algorithms, that have one thing in common: they simulate a Wiener process on a dataset and directly use some properties of time-evolving density to estimate LID. We can divide those algorithms into the following three groups:
@@
    LIDL (Tempczyk et al. 2022) and its efficient and most accurate implementation using diffusion models called FLIPD (Kamkari et al. 2024b). These algorithms use the rate of change of the probability density at point $x$ during the Wiener process to estimate LID at $x$. For small diffusion times $t$ the logarithm of a density is a linear function of a logarithm of $t$, and the proportionality coefficient is equal $d-D$ for small $t$, where $d$ is manifold density and $D$ is ambient space dimensionality. Our experiments with ID-NF (described below) and FLIPD show, that the latter is more scalable than ID-NF (and ID-DM) due to the high memory and computational complexity of SVD (which has to be calculated for each data point). Experiments from (Kamkari et al. 2024b) show that FLIPD is more accurate than NB (described below), which led us to the conclusion, that FLIPD is state-of-the-art in LID estimation among the most scalable algorithms.
@@
    ID-NF (Horvat and Pfister 2022) (using normalizing flows), and the diffusion-using follow-up paper ID-DM (Horvat and Pfister 2024) analyze how singular values of a Jacobian of a function transforming a standard normal distribution into a diffused dataset density at $x$ evolves during the Wiener process. Horvat and Pfister observed that when transforming a $d$-dimensional manifold into a $D$-dimensional Gaussian we have to expand space more in the normal direction to manifold, especially for small diffusion times $t$.
@@
    NB (Stanczuk et al. 2024), which is an abbreviation from Normal Bundle (name used in (Kamkari et al. 2024b)). Stanczuk et al. observed that for small diffusion times $t$ the gradients of the logarithm of a diffused data density (score function) close to $x$ lies in the normal space to the manifold and use this fact to estimate LID at $x$.
@@ 3 Isolated and holistic algorithms
In this work, we take a closer look at the Wiener process-based algorithms described in the last paragraph of Section 2. Although all these algorithms apply a Wiener process to the dataset during their first phase, when looking at their second step, we can divide them into two groups: isolated (LIDL, FLIPD, NB) and holistic (ID-NF, ID-DM). The intermediate results of the first group that are used for LID calculation use only the information about the local shape of the data probability density function (without normalization constant). We assume that the generative model approximates diffused data distribution $\rho_{t}$ perfectly. Their estimates depend on the proximity of $\nabla\log\rho_{t}$ to $x$ (NB) or $d\log\rho_{t}/d\log t$ at $x$ (LIDL, FLIPD).
This is a consequence of $\rho_{t}$ at $x$ being a function that – in practice – depends on the values of original data distribution $p_{S}$ in a ball of radius $r\approx 4\sqrt{t}$ around $x$ in the data space. The values of $p_{S}$ outside this ball do not matter in practice for isolated algorithms, because the diffused particles in the Wiener process can travel longer distances than $r$ with very low probability (less than $3.7\cdot 10^{-5}$). When we add some new data points to the dataset far away from $x$, it does not change the shape of $\rho_{t}$ close to $x$. This operation only changes a normalization constant, which becomes 0 after taking a logarithm of $\rho_{t}$ and taking a derivative either w.r.t time or spatial variables.
The holistic algorithms work quite differently. In the case of ID-NF and ID-DM, they calculate singular values of the Jacobian of the function $\zeta:\mathrm{R}^{N}\rightarrow\mathrm{R}^{N}$ transforming $\rho_{t}$ into a standard normal distribution. This $\zeta$ function strongly depends on the entire shape of $\rho_{t}$. As mentioned before, when we change $\rho_{t}$ far away from $x$ we do not change the shape of $\rho_{t}$ close to $x$. The same is not true for $\zeta(x)$. When we add many data points to the dataset far away from $x$ while we keep the latent distribution fixed, we have to change the way we compress and stretch the space to match the new distribution. This property makes the analysis of the behavior of ID-NF and ID-DM much harder (maybe even impossible) if we want to take into account only the data density in the neighbourhood of $x$.
To give an illustrative example of this behavior, one needs to imagine that our dataset is one-dimensional and consists of 10K points sampled from $\mathcal{N}\left(0,1\right)$. Typically, we are transforming it into $\mathcal{N}\left(0,1\right)$, so $\zeta$ is just an identity function. Now let’s add to the dataset another 10K points sampled from $\mathcal{N}\left(100,1\right)$. As a consequence, we have to stretch our space in some areas to transform one density into another, whereas $\zeta$ changes along with its Jacobian.
@@ 4 Wiener process perspective on LID estimation
In this section we present a new perspective on perturbing datasets, unifying the approaches seen in the algorithms presented by Tempczyk et al. (2022); Stanczuk et al. (2024); Horvat and Pfister (2022, 2024); Kamkari et al. (2024b). As already mentioned, all these algorithms consist of two stages, the first of which amounts to perturbing the dataset with normally distributed random noise of fixed variance $t$. In the second stage, each of the algorithms utilizes the behavior of the perturbed density in the neighborhood of a fixed point under changes in the noise variance.
The first phase of each algorithm can be interpreted as applying the Wiener process to the points in the dataset. Afterward, the resulting set of points is used to train some type of generative model (or models) to estimate the distribution of the dataset undergoing the Wiener process at time $t$. From the point of view of differential equations, the distribution density function of the diffused dataset is described by Fick’s Second Law of Diffusion.
@@
Now, given a dataset embedded in $\mathbb{R}^{D}$, we assume that it has been drawn from some latent union of submanifolds $S$ endowed with a probability measure $p_{S}$ (which can be naturally treated as a probability measure on $\mathbb{R}^{D}$). The goal of Local Intrinsic Dimension estimation is to find out the dimension of $S$ at any point of the dataset.
@@
This is the density of normal distribution on $\mathbb{R}^{D}$ with covariance matrix $tI$. It is the fundamental solution of the differential equation given by Fick’s Second Law of Diffusion (1) with $C=1/2$. Here, this means that the convolution
$\rho_{t}=p_{S}*\phi_{t}^{D}$ (3)
@@
To limit the complexity introduced by curvature, from now on we will consider only flat manifolds. This means that, without loss of generality, we may assume that $S$ is the first factor in product decomposition $\mathbb{R}^{D}=\mathbb{R}^{d}\times\mathbb{R}^{D-d}$. We will denote the coordinates of $\mathbb{R}^{d}$ and $\mathbb{R}^{D-d}$ by $x$ and $y$, respectively. We will moreover assume that $p_{S}$, now a probability distribution on $\mathbb{R}^{d}$, has a density $\psi\colon\mathbb{R}^{d}\to\mathbb{R}$.
In Appendix B we discuss the Laplacian of $\rho_{t}$ and derive the following result:
Lemma 4.1 (Lemma A Wiener Process Perspective on Local Intrinsic Dimension Estimation Methods).
For $t>0$ and $(x,y)\in\mathbb{R}^{D}$ we have
$\begin{split}\Delta\rho_{t}(x,y)&=\left(\frac{\left\lVert{y}\right\rVert^{2}}{t^{2}}+\frac{d-D}{t}\right)\rho_{t}(x,y)\\ &+\phi_{t}^{D-d}(y)\Delta_{x}(\psi*\phi_{t}^{d})(x).\end{split}$ (4)
As a consequence, by putting $y=0$ and using
$\phi_{t}^{D-d}(0)=(2\pi t)^{(d-D)/2}$ (5)
we obtain the following.
Corollary 4.2.